OpenStudy (anonymous):

Can you please help me with this? This is part of an assignment that is due tonight.

OpenStudy (anonymous):

genson0 If F is the focal length of a convex lens and an object is placed at a distance x from the lens, then its image will be at a distance y from the lens, where F, x, and y are related by the lens equation \[\frac{ 1 }{ F }=\frac{ 1 }{ x }+\frac{ 1 }{ y }\] Suppose that a lens has a focal length of 4.2 cm, and that the image of an object is 8 cm closer to the lens than the object itself. How far from the lens is the object?

OpenStudy (anonymous):

Could you explain please? Sorry.

OpenStudy (anonymous):

6.1cm doesn't seem to be write according to webassign.

OpenStudy (anonymous):

the question is, what is "x" ? the information given is F= 4.2 and " image of an object is 8 cm closer to the lens than the object itself" which can be expressed as the equation y= x-8

OpenStudy (anonymous):

substituting the given information into the given equation \[\frac{1}{4.2}= \frac{1}{x} + \frac{1}{x-8}\] from here, you just need to solve for x

OpenStudy (anonymous):

Yes I got to this point, I am lost at this point. Please help.

OpenStudy (anonymous):

combine the right side

OpenStudy (anonymous):

what is the common denominator?

OpenStudy (anonymous):

I am at this point \[\frac{ d(d-8) }{ 4.2d(d-8) }=\frac{ 4.2d -33.6 +4.2d }{ 4.2d(d-8 }\]

OpenStudy (anonymous):

right, now get rid of the denominators

OpenStudy (anonymous):

you if multiplied 4.2 d ( d- 8) to both sides, then the denominators would cancel out

OpenStudy (anonymous):

then move everything to one side, at this point, you will have a 2nd degree polynomial equal to zero solve either by factoring or by using the quadratic equation

OpenStudy (anonymous):

\[d^2-16.4d+33.6=0\] Alright this is where I dont know, I dont know any factor of 16.4 and 33.6

OpenStudy (anonymous):

Are these the right numbers?

OpenStudy (anonymous):

i have no idea, i dont have a piece of paper in front of me

OpenStudy (anonymous):

assuming you didnt make a mistake earlier, the numbers seem fine

OpenStudy (anonymous):

I'm lost at this point :/ I don't know what numbers to instert

OpenStudy (anonymous):

quadratic equation for \[ax^2+bx+c = 0\] \[x= \frac{-b\pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

OH how could i not think of that.

OpenStudy (anonymous):

I got d=14, and d=2.4 So how do I get the answer? it asks for only one

OpenStudy (anonymous):

They both work also into the equation.

OpenStudy (anonymous):

y= x-8 which answer makes sense now?